[Hallc_running] [New Logentry] Status update

sanghwa at jlab.org sanghwa at jlab.org
Sat Apr 7 10:55:03 EDT 2018

Logentry Text:

Yesterday the MCC did pass change for Hall B (3->5) started around 2:30pm. During the downtime, we had a short escorted access. Brad went down to power cycle HCCAM01(SHMS side), but the camera is dead (https://logbooks.jlab.org/entry/3556053). We will need to live with the other two cameras for awhile.

In the meantime, we changed the HMS setting to Pm = 750 MeV setting for the D2 Electro-Disintegration Run. We started collecting data at the Pm = 750 MeV setting since the swing shift yesterday. We have collected ~46% of the minimum charge goal (5270 mC) and 23% of the actual goal (10540) so far for this kinematic setting. MCC kept delivering 60uA to Hall C. There was a 30min downtime around 5:20 am due to a vacuum issue, but otherwise it has been smooth data taking.

The plan is to continue taking data over the weekend. There will be a scheduled Beam Study (8hrs) on Monday starting from 8:30.

There will be no RC meeting today. Have a great weekend everyone and wish us happy data taking!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3556468
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180407/b787509a/attachment.html>

More information about the Hallc_running mailing list